# A defect-chemistry-informed phase-field model of grain growth in oxide electroceramics

### v2

@@ 3.1 Free energy density functional and electrochemical potentials
$\begin{split}f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})&=\frac{1}{V_{\text{m}}}\left\{g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}C_{\text{V}_{\text{O}}^{..},\text{b}}+g^{0}_{\text{dop},\text{b}}C_{\text{dop},\text{b}}\right.\\ &+RT\left[C_{\text{V}_{\text{O}}^{..},\text{b}}\ln(C_{\text{V}_{\text{O}}^{..},\text{b}})+(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}})\ln(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}})\right.\\ &\left.\left.+C_{\text{dop},\text{b}}\ln(C_{\text{dop},\text{b}})+(\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}})\ln(\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}})\right]\right\}\\ &+\frac{\mathcal{F}}{V_{\text{m}}}\left(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..},\text{b}}+z_{\text{dop}}C_{\text{dop},\text{b}}\right)\phi,\end{split}$ (27)
where $g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}$, $g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}$, $g^{0}_{\text{dop},\text{c}}$, $g^{0}_{\text{dop},\text{b}}$ are standard formation energies of oxygen vacancy and dopant in bulk and core. $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$, $\tilde{N}_{\text{dop},\text{c}}$, $\tilde{N}_{\text{dop},\text{b}}$ are dimensionless numbers of available sites per unit volume for oxygen vacancy and dopant in bulk and core. $z_{\text{V}_{\text{O}}^{..}}$ and $z_{\text{dop}}$ are valance numbers of oxygen vacancy and dopant. $\mathcal{F}$, $V_{\text{m}}$, $R$, $T$ are the Faraday constant, molar volume, gas constant and temperature, respectively. The electrochemical potentials of oxygen vacancy and dopant in bulk and core are obtained by
@@
$\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{b}}}=\frac{g_{\text{V}_{\text{O}}^{..},\text{b}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{V}_{\text{O}}^{..},\text{b}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{V}_{\text{O}}^{..}}\phi,$ (28)
@@
$\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{c}}}=\frac{g_{\text{V}_{\text{O}}^{..},\text{c}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{V}_{\text{O}}^{..},\text{c}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{V}_{\text{O}}^{..}}\phi,$ (29)
@@
$\mu_{\text{dop},\text{b}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{b}}}=\frac{g_{\text{dop},\text{b}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{dop},\text{b}}}{\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{dop}}\phi,$ (30)
@@
$\mu_{\text{dop},\text{c}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{c}}}=\frac{g_{\text{dop},\text{c}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{dop},\text{c}}}{\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{dop}}\phi,$ (31)
@@ 3.2 Governing equations
In the framework of KKS phase-field model, the evolution of phase-field OPs obeys the Allen-Cahn equation. For grain $i$, the evolution of $\eta_{i}$ in time is
@@
$\begin{split}\frac{v_{c}}{L}\frac{d\eta}{dx}&=\nabla\cdot\kappa\nabla\eta_{i}-\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\\ &+\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right],\end{split}$ (34)
Conserved concentration fields $C_{\text{V}_{\text{O}}^{..}}$ and $C_{\text{dop}}$ follow Fick’s law, which are expressed as
@@
Therefore, $\delta\mathscr{F}/\delta C_{\text{V}_{\text{O}}^{..}}$ and $\delta\mathscr{F}/\delta C_{\text{dop}}$ in Eqs. 35 and 36 can be substituted by $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{\text{PF}}$, $\mu_{\text{dop},\text{b}}^{\text{PF}}$ or $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{\text{PF}}$, $\mu_{\text{dop},\text{c}}^{\text{PF}}$. In addition, the governing equation of electrostatic potential field is obtained by Poisson’s equation
@@
$\epsilon_{0}\epsilon_{\text{r}}\nabla^{2}\phi+\frac{\mathcal{F}}{V_{\text{m}}}(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..}}+z_{\text{dop}}C_{\text{dop}})=0,$ (39)
The parameters utilized in phase-field model are tabulated in table 2.

### v3

@@ 3.1 Free energy density functional and electrochemical potentials
$\begin{split}f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})&=\frac{1}{V_{\text{m}}}\left\{g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}C_{\text{V}_{\text{O}}^{..},\text{b}}+g^{0}_{\text{dop},\text{b}}C_{\text{dop},\text{b}}\right.\\ &+RT\left[C_{\text{V}_{\text{O}}^{..},\text{b}}\ln\left(\frac{C_{\text{V}_{\text{O}}^{..},\text{b}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}}\right)+(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}})\ln\left(\frac{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}}\right)\right.\\ &\left.\left.+C_{\text{dop},\text{b}}\ln\left(\frac{C_{\text{dop},\text{b}}}{\tilde{N}_{\text{dop},\text{b}}}\right)+(\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}})\ln\left(\frac{\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}}}{\tilde{N}_{\text{dop},\text{b}}}\right)\right]\right\}\\ &+\frac{\mathcal{F}}{V_{\text{m}}}\left(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..},\text{b}}+z_{\text{dop}}C_{\text{dop},\text{b}}\right)\phi,\end{split}$ (28)
where $g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}$, $g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}$, $g^{0}_{\text{dop},\text{c}}$, $g^{0}_{\text{dop},\text{b}}$ are standard formation energies of oxygen vacancy and dopant in bulk and core. $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$, $\tilde{N}_{\text{dop},\text{c}}$, $\tilde{N}_{\text{dop},\text{b}}$ are dimensionless numbers of available sites per unit volume for oxygen vacancy and dopant in bulk and core. $z_{\text{V}_{\text{O}}^{..}}$ and $z_{\text{dop}}$ are valance numbers of oxygen vacancy and dopant. $\mathcal{F}$, $V_{\text{m}}$, $R$, $T$ are the Faraday constant, molar volume, gas constant and temperature, respectively. Based on experimental evidence of abnormal grain growth in 0.2% Fe-doped STO zahler2023grain , where defect-defect interactions are negligible, we limit our consideration to the dilute defect case. This simplification ensures that the fundamental characteristics of the system are captured without introducing unnecessary complexities. However, the dilute assumption is not valid when high defect concentration is involved, the interactions between different defects should be considered chen2022thermodynamic .
The electrochemical potentials of oxygen vacancy and dopant in bulk and core are obtained according to
@@
$\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{b}}}=\frac{g_{\text{V}_{\text{O}}^{..},\text{b}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left(\frac{C_{\text{V}_{\text{O}}^{..},\text{b}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}}}\right)}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{V}_{\text{O}}^{..}}\phi,$ (29)
@@
$\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{c}}}=\frac{g_{\text{V}_{\text{O}}^{..},\text{c}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left(\frac{C_{\text{V}_{\text{O}}^{..},\text{c}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}}}\right)}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{V}_{\text{O}}^{..}}\phi,$ (30)
@@
$\mu_{\text{dop},\text{b}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{b}}}=\frac{g_{\text{dop},\text{b}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left(\frac{C_{\text{dop},\text{b}}}{\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}}}\right)}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{dop}}\phi,$ (31)
@@
$\mu_{\text{dop},\text{c}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{c}}}=\frac{g_{\text{dop},\text{c}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left(\frac{C_{\text{dop},\text{c}}}{\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}}}\right)}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{dop}}\phi,$ (32)
@@ 3.2 Governing equations
As mentioned in the introduction, we focus on the solute drag effect, which is the most fundamental and primary factor influencing grain growth patterns. Therefore, our phase-field model employs constant GB energy and mobility and uses order parameters to distinguish different grains without explicitly considering grain misorientation. Then, the evolution of phase-field OPs for grain $i$ with isotropic GB properties becomes
@@
$\begin{split}\frac{v_{c}}{L}\frac{d\eta}{dx}&=\nabla\cdot\kappa\nabla\eta_{i}-\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\\ &+\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right].\end{split}$ (35)
By combining Fick’s first and second laws, the evolution of the conserved concentration fields $C_{\text{V}_{\text{O}}^{..}}$ and $C_{\text{dop}}$ are expressed as
@@
Therefore, $\delta\mathscr{F}/\delta C_{\text{V}_{\text{O}}^{..}}$ and $\delta\mathscr{F}/\delta C_{\text{dop}}$ in Eqs. 36 and 37 can be substituted by $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{\text{PF}}$, $\mu_{\text{dop},\text{b}}^{\text{PF}}$ or $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{\text{PF}}$, $\mu_{\text{dop},\text{c}}^{\text{PF}}$. In addition, the governing equation of electrostatic potential field is obtained by Poisson’s equation
[…]
Substituting the site fractions $_{\text{V}_{\text{O}}^{..},\text{b}}], [VV$$_{\text{V}_{\text{O}}^{..},\text{c}}$], $V_{\text{dop},\text{b}}]$ and $V_{\text{dop},\text{c}}]$ into Eq. 40, we have
@@
$\begin{split}\epsilon_{0}\epsilon_{\text{r}}\nabla^{2}\phi=&-\frac{\mathcal{F}}{V_{\text{m}}}\left\{z_{\text{V}_{\text{O}}^{..}}\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}[V_{\text{V}_{\text{O}}^{..},\text{b}}][1-h(\eta)]+z_{\text{V}_{\text{O}}^{..}}\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}[V_{\text{V}_{\text{O}}^{..},\text{c}}]h(\eta)\right.\\ &\left.+z_{\text{dop}}\tilde{N}_{\text{dop},\text{b}}[V_{\text{dop},\text{b}}][1-h(\eta)]+z_{\text{dop}}\tilde{N}_{\text{dop},\text{c}}[V_{\text{dop},\text{c}}]h(\eta)\right\}.\end{split}$ (41)
We clearly observe the numbers of available sites $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$, $\tilde{N}_{\text{dop},\text{b}}$, $\tilde{N}_{\text{dop},\text{c}}$ influence the electrostatic potential distribution in Eq. 41. The parameters utilized in phase-field model are tabulated in table 2.
3.3 Comparison to other grain growth phase-field models for oxide ceramics
In this part, we outline differences between our defect-chemistry-informed phase-field model, the model developed by Vikrant et al. vikrant2020electrochemical and the grand-potential based phase-field model developed by Aagesen et al. aagesen2024electrochemical .
In accordance with defect chemistry theory maier2023physical , we include the available site densities for the bulk and the grain boundary core explicitly, as shown by the bulk and GB core electrochemical free energy (Eq. 27 and Eq. 28) and the oxygen vacancy and dopant electrochemical potentials in bulk and GB core (Eq. 29-Eq. 32). It is important to do so, because they can impact the space charge layers and GB potential considerably usler2024point . As mentioned in the Introduction, in Fe-doped SrTiO3, Fe substitutes for Ti4+ on the B-site, and charge compensation is typically provided by oxygen vacancies on the O-site. This results in the number density of available sites for the dopant in the bulk that is one-third of that for an oxygen vacancy. Even more important, B-site dopant and O-site oxygen vacancy belong to different sub-lattices and thus have to be accounted for separately, in order to correctly represent the respective configurational entropy. In addition, the number of available sites for both species will deviate significantly from their bulk values due to the distinct atomic structure of the GB core. Therefore, it is necessary to treat the available site densities for oxygen vacancies and dopants in the GB core as distinct from those in the bulk, and these quantities play a critical role in determining the space charge layer. As our simulation results in the following sections demonstrate, the different numbers of available site significantly influence both the grain boundary potential and the defect concentration distributions, consequently, leading to different space charge layers formation. Much more results can be found in Figures 2 and 3, validated against De Souza’s work de2009formation .
Such explicit incorporation of the numbers of available site of oxygen vacancy and dopant in the bulk and GB core is not fully regarded in the existing models. For instance, Vikrant et al. assumed that the oxygen vacancy and acceptor dopant share the same sublattice in the chemical free energy contribution (see Eq. 1 in Ref. vikrant2020electrochemical ).
$\begin{split}f^{\text{Vik}}(\eta,[V_{\text{V}_{\text{O}}^{..}}],[V_{\text{dop}}],T)&=\frac{1}{\omega}\left\{f_{\text{V}_{\text{O}}^{..}}(\eta,T)[V_{\text{V}_{\text{O}}^{..}}]+f_{\text{dop}}(\eta,T)[V_{\text{dop}}]+k_{\text{B}}T\left([V_{\text{V}_{\text{O}}^{..}}]\ln[V_{\text{V}_{\text{O}}^{..}}]+[V_{\text{dop}}]\ln[V_{\text{dop}}]\right)\right.\\ &\left.+k_{\text{B}}T\left(1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]\right)\ln\left(1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]\right)+\Omega_{\text{V}_{\text{O}}^{..}\text{dop}}[V_{\text{V}_{\text{O}}^{..}}][V_{\text{dop}}]\right\},\end{split}$ (42)
where $f_{\text{V}_{\text{O}}^{..}}(\eta,T)$ and $f_{\text{dop}}(\eta,T)$ are temperature related formation energy of oxygen vacancy and dopant in bulk and GB core. $V_{\text{V}_{\text{O}}^{..}}]$ and $V_{\text{dop}}]$ are the site fractions of oxygen vacancy and dopant. $\Omega_{\text{V}_{\text{O}}^{..}\text{dop}}$ is the interaction coefficient between oxygen vacancy and dopant. At here, we omit $\Omega_{\text{V}_{\text{O}}^{..}\text{dop}}$ for dilute case. In Eq. 42, both oxygen vacancy and dopant are described as occupying the same sublattice, which is reflected in the entropic mixing term $\left(1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]\right)\allowbreak\ln\left(1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]\right)$. Then, we can obtain the electrochemical potential of oxygen vacancy and dopant from Vikrant’s model
$\mu_{\text{V}_{\text{O}}^{..}}^{\text{Vik}}=\frac{f_{\text{V}_{\text{O}}^{..}}(\eta,T)}{\omega}+\frac{k_{\text{B}}T}{{\omega}}\ln\left(\frac{[V_{\text{V}_{\text{O}}^{..}}]}{1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]}\right)+z_{\text{V}_{\text{O}}^{..}}e\phi,$ (43)
$\mu_{\text{dop}}^{\text{Vik}}=\frac{f_{\text{dop}}(\eta,T)}{\omega}+\frac{k_{\text{B}}T}{\omega}\ln\left(\frac{[V_{\text{dop}}]}{1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]}\right)+z_{\text{dop}}e\phi,$ (44)
which is different from the electrochemical potential obtained from defect-chemistry as shown in Eq. 1-Eq. 5 and present work in Eq. 29 - Eq. 32. In contrast, our model explicitly distinguishes between the numbers of available site of oxygen vacancy and dopant in both bulk and GB core, i.e. $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$, $\tilde{N}_{\text{dop},\text{b}}$, $\tilde{N}_{\text{dop},\text{c}}$. The configurational entropy term in the electrochemical free energy part is $\left(\tilde{N}_{i}-C_{i}\right)\ln\left(\frac{\tilde{N}_{i}-C_{i}}{\tilde{N}_{i}}\right)$, with the subscript $i$ indicates $\text{V}_{\text{O}}^{..},\text{b}$, $\text{V}_{\text{O}}^{..},\text{c}$, $\text{dop},\text{b}$ and $\text{dop},\text{c}$. This separation allows us to represent the thermodynamics of defect segregation more accurately at grain boundary core. Moreover, the numbers of available site of oxygen vacancy and acceptor dopant in both the bulk and GB core explicitly influence the Poisson’s equation and the electrostatic potential distribution in our model (see Eq. 41), However, in Vikrant’s model, the Poisson’s equation is $\epsilon_{0}\epsilon_{r}\nabla^{2}\phi=-e(z_{\text{V}_{\text{O}}^{..}}[V_{\text{V}_{\text{O}}^{..}}]+z_{\text{dop}}[V_{\text{dop}}])$, the influence of these parameters, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$, $\tilde{N}_{\text{dop},\text{b}}$, $\tilde{N}_{\text{dop},\text{c}}$, on the electrostatic potential distribution is not considered.
Aagsen’s model is developed based on the grand potential functional. The thermodynamic equivalence between the free energy based KKS model and grand potential based phase-field model developed has been clarified in Ref. plapp2011unified (see section VA). However, in Aagsen’s model, the numbers of available site (denoted as $n_{\text{Y}}$ and $n_{\text{O}}$) of both cation vacancy and oxygen vacancy in Y2O3 are treated as uniform across the bulk and GB core, i.e. $n_{\text{Y,b}}=n_{\text{Y,c}}$ and $n_{\text{O,b}}=n_{\text{O,c}}$, which may overlook the important structural variations at GB core, especially when $n_{\text{Y,b}}\neq n_{\text{Y,c}}$ and $n_{\text{O,b}}\neq n_{\text{O,c}}$, and result in inappropriate space charge layer formation. These scenarios have been comprehensively discussed in De Souza’s work de2009formation and demonstrated in the present work (see Fig. 2 and Fig. 3) in Section 5.1.
In addition, such explicit consideration of numbers of available site in our formulation paves the foundation for the subsequent extension towards the treatment of charge state transition of dopants, such as Fe4+/Fe3+ and Fe3+/Fe2+ transitions in Fe-doped SrTiO3 klein2023fermi, bonkowski2024single. Vikrant’s work, Aagsen’s work and the present work primarily focus on a single valence state (typically -1), which limits their applicability under varying thermodynamic conditions. The dominant charge state of Fe can vary significantly with temperature and oxygen partial pressure wang2016defect . These charge states all share the same B-site sublattice in the perovskite structure, which reinforces the importance of correctly defining and distinguishing the numbers of available site between oxygen vacancy and dopant in bulk phase and grain boundary core in the phase-field model. Without completely considering the defect chemistry, the extension of Vikrant’s model or Aagsen’s model in this direction is not applicable due to the mixture of the numbers of available site. However, our defect-chemistry phase-field is developed to accommodate this level of resolution, enabling us to simulate space charge layer formation with proper consideration of the different valence states of dopant and their impact on different space charge layer formations.